OpenStudy (pinklion23):

32-6x+7x=(-15)+28 solve for x X=

jagr2713 (jagr2713):

Hey what do you think we do first?

OpenStudy (pinklion23):

subtract 32 from both sides

jagr2713 (jagr2713):

Yes you can but don't you want to add like terms first?

OpenStudy (pinklion23):

oh yeah i forget so 13x

jagr2713 (jagr2713):

you forgot the negative sign.

OpenStudy (pinklion23):

oh ok so 32-(-13x)= (-15)+28?

jagr2713 (jagr2713):

nope -6+7x= 1

OpenStudy (pinklion23):

so then x=1

jagr2713 (jagr2713):

Yes so we have 32+x=(-15)+28 left right? so what next

OpenStudy (pinklion23):

then we do32- 28 so then it would be 4+x=(-15)

jagr2713 (jagr2713):

good what next?

OpenStudy (pinklion23):

the x=-19?

jagr2713 (jagr2713):

Correct :D

OpenStudy (pinklion23):

wow thx so much can you help me with some more?

jagr2713 (jagr2713):

Of course :D

OpenStudy (pinklion23):

-4X+5(X-2)=-16+10

jagr2713 (jagr2713):

what do we do first?

OpenStudy (pinklion23):

-16+10?

jagr2713 (jagr2713):

yea you can so we get -6

jagr2713 (jagr2713):

-4x+5(x-2)=-6 whats next?

OpenStudy (pinklion23):

yeah

OpenStudy (pinklion23):

um then we do -6-2?

jagr2713 (jagr2713):

Nope when we have a number like 5(x-2) we distribute. Do you how to distribute ?

OpenStudy (pinklion23):

so then we do (x*5)-(2*5)

jagr2713 (jagr2713):

Yes like this

jagr2713 (jagr2713):

So we just multiply everything inside that parenthesis

OpenStudy (pinklion23):

yeah

jagr2713 (jagr2713):

so now we got -4x+5x-10=-6 whats next

OpenStudy (pinklion23):

then we have 1x-10=-6

jagr2713 (jagr2713):

It's not 1x x=1 so u don't need to put the 1

OpenStudy (pinklion23):

the we add -10 to -6

OpenStudy (pinklion23):

?

OpenStudy (pinklion23):

so X=1

jagr2713 (jagr2713):

you got that? x=1 you out 1x-10=-6 but x already is 1 so u don't need to put the 1x just x-10=-6

OpenStudy (pinklion23):

ok so X = -16

OpenStudy (pinklion23):

or -4

jagr2713 (jagr2713):

nope -6+10?

jagr2713 (jagr2713):

positive 4 :D

OpenStudy (pinklion23):

oh ok thx

jagr2713 (jagr2713):

Yep Anytime :D

OpenStudy (pinklion23):

i have two more can you help me with them

jagr2713 (jagr2713):

ummm ill see what i can do :D